OpenStudy (anonymous):

how many 3 letter permutations are possible for the letters s,t,u,d,y,h,a,r,d

OpenStudy (anonymous):

there are 14 :)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

i have another question

OpenStudy (anonymous):

bella

OpenStudy (anonymous):

Sure what is it?

OpenStudy (anonymous):

jason and kyle both choose a number from 1 to 10 at random. what is the probability that both numbers are odd?

OpenStudy (anonymous):

Jason and Kyle both choose a number from 1 to 10 at random. What is the probability that both numbers are odd? --- Jason chooses odd: 5/10 = 1/2 Kyle chooses odd: 5/10 = 1/2 P(both choose odd) = (1/2)^2 = 1/4

OpenStudy (cwrw238):

@bella - it would be better for randilauren if you explained how you obtained the answer to the first question

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

@cwrw238 can you explain it? I have to go help my mom :)